# Banana Trees for the Persistence in Time Series Experimentally

### v1

@@ C.2.2 Additional Results
Figure 12: Timing the maintenance of banana trees relative to Gudhi on linear-time inputs with varying noise parameter $\lambda$. The baseline of no speedup ($110^{0}=$ 1) is shown as a horizontal gray line. 12 Median speedup for large updates in the linear-time case for local maintenance. 12 Median speedup in the linear-time case for cutting. The peaks in the bottom right for $\lambda=0$ (yellow line) originate in the running time for Gudhi and are not observed in the time for banana trees. 12 Median speedup in the linear-time case for gluing.
@@
C.3 Quasi-Periodic Inputs
@@
Figure 13: Timing the construction of banana trees for quasi-periodic inputs. Both plots show the slowdown over Gudhi. Ribbons extend from the minimum to the maximum observed value; dots show the mean. The baseline of no slowdown is shown as a horizontal gray line. 13 Slowdown for $\sigma=1$ dependent on input size $n$. 13 Slowdown dependent on $\sigma$ and averaged over all input sizes.
@@
Figure 14: Speedup of banana trees relative to Gudhi when performing local and topological updates on quasi-periodic inputs. The baseline of no speedup is shown as a horizontal gray line. 14 Speedup when performing local updates to inputs of size $n$ and with parameter $\sigma=1$. The ribbon extends from the minimum to the maximum observed value; the line traces the median speedup. Banana trees are faster than Gudhi on all inputs with $n\geq 1683$; see vertical dashed line. In the median they are faster on all inputs with $n\geq 264$, as shown by the dotted line. 14 Speedup of cutting inputs of size $n$ with parameter $\sigma$ in half. 14 Speedup of gluing two equally large inputs with parameter $\sigma$ into a list of size $n$.
@@
C.4 Real-World Examples

### v2

@@ B.2.2 Additional Results
Figure 13: Timing the maintenance of banana trees relative to Gudhi on linear-time inputs with varying noise parameter $\lambda$. The baseline of no speedup ($110^{0}=$ 1) is shown as a horizontal gray line. 13 Median speedup for large updates in the linear-time case for local maintenance. 13 Median speedup in the linear-time case for cutting. The peaks in the bottom right for $\lambda=0$ (yellow line) originate in the running time for Gudhi and are not observed in the time for banana trees. 13 Median speedup in the linear-time case for gluing.
@@
B.3 Quasi-Periodic Inputs
@@
Figure 14: Timing the construction of banana trees for quasi-periodic inputs. Both plots show the slowdown over Gudhi. Ribbons extend from the minimum to the maximum observed value; dots show the mean. The baseline of no slowdown is shown as a horizontal gray line. 14 Slowdown for $\sigma=1$ dependent on input size $n$. 14 Slowdown dependent on $\sigma$ and averaged over all input sizes.
@@
Figure 15: Speedup of banana trees relative to Gudhi when performing local and topological updates on quasi-periodic inputs. The baseline of no speedup is shown as a horizontal gray line. 15 Speedup when performing local updates to inputs of size $n$ and with parameter $\sigma=1$. The ribbon extends from the minimum to the maximum observed value; the line traces the median speedup. Banana trees are faster than Gudhi on all inputs with $n\geq 1683$; see vertical dashed line. In the median they are faster on all inputs with $n\geq 264$, as shown by the dotted line. 15 Speedup of cutting inputs of size $n$ with parameter $\sigma$ in half. 15 Speedup of gluing two equally large inputs with parameter $\sigma$ into a list of size $n$.
@@
B.4 Real-World Examples